OpenStudy (anonymous):

I'm trying to sketch a curve can someone check what I have so far: Domain: all real numbers except 8 y intercept: 8 x intercept: none Vertical A: 8 horizontal A: 0 function: f(x)= x-8/x^2 first derivative: 1+16/x^3 critcal number = 8 second derivative: -48/x^4

OpenStudy (anonymous):

try facebook maths nerds group they anser ur questions

OpenStudy (anonymous):

Is the function: \(\displaystyle f(x)=x-\frac{8}{x^2}\) or \(\displaystyle f(x)=\frac{x-8}{x^2}\) Also the asymptotes and intercepts are wrong

OpenStudy (anonymous):

\[f(x)= x-8/x^2\]

OpenStudy (anonymous):

The function explodes at zero. So you will have a vertical asymptote there. You will also have an oblique asymptote because the limiting behaviour of the function is \(g(x)=x\)

OpenStudy (anonymous):

\[\lim_{x \to \infty} x-\frac{8}{x^2} = x\]

OpenStudy (anonymous):

I take it thats the horixonal asymptote right?

OpenStudy (anonymous):

It's an oblique asymptote, not a horizontal asymptote. http://en.wikipedia.org/wiki/Asymptote#Oblique_asymptotes

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

You are considering what happens when x gets very large (negative or positive). Think about what happens.

OpenStudy (anonymous):

The term \(\displaystyle\frac{8}{x^2}\) collapses

OpenStudy (anonymous):

All you are left with is \(x\)

OpenStudy (anonymous):

oh ok so for example this function: \[f(x)=x^2+1/x\] would have a slant asmpotote as well because its undefined at 0?

OpenStudy (anonymous):

You will have a vertical asymptote at zero, since the term \(\displaystyle\frac{1}{x}\) explodes at 0. You will have a quadratic asymptote for large x.